OpenStudy (anonymous):

HELP ME WITH THIS 2x-1<=5<=x+2 i have to find the interval and plot it on a number line

OpenStudy (anonymous):

2x-1=<5 or 5=<x+2. (split them up) 2x=<6. -x=<2-5 x=<3. -x=<-3 x>=3 Thus x=3 only

OpenStudy (jdoe0001):

\(\large{ 2x-1\le 5\le x+2\implies \begin{cases} 2x-1\le 5\\ \quad \\ 5\le x+2 \end{cases}\\\quad \\ \begin{cases} 2x-1\le 5\implies x \le \cfrac{6}{2}\implies &x\le 3\\ \quad \\ 5\le x+2\implies &3\le x \end{cases}\implies x =3}\)

OpenStudy (anonymous):

thank you guys :)